OpenStudy (anonymous):

A = [5,8/3,-2/3; 2,2/3,4/3; -4, -4/3, -8/3] (Matrix form 3x3) How do I write down a matrix that diagonalises A? Please show all working.

OpenStudy (zarkon):

find the eigenvectors

OpenStudy (anonymous):

yep i have got them whats next?

OpenStudy (zarkon):

form a matrix P then \[P^{-1}AP\] is a diagonal matrix

OpenStudy (anonymous):

so basically form a 3x3 out of the eigenvectors

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Zarkon would you mind giving this a go nd posting your solutions?

OpenStudy (anonymous):

i know the eigenvalues are: lambda = 0 lambda = -3 lambda = 6

OpenStudy (zarkon):

correct

OpenStudy (zarkon):

I get \[P=\left[\begin{matrix}-2 & 1/4 & -2 \\ -1/2 & -1/2 & 4\\ 1 & 1& 1\end{matrix}\right]\] then \[P^{-1}AP=\left[\begin{matrix}6 & 0 & 0 \\ 0 & -3 & 0\\ 0 & 0& 0\end{matrix}\right]\]

OpenStudy (anonymous):

yes that is correct although what are the workings for it.

OpenStudy (zarkon):

Compute the null space for the matrices \[A-\lambda I\] \[\text{Where }\lambda=6,-3,0\]

OpenStudy (anonymous):

ok thanks i will keep going you are too good.

OpenStudy (anonymous):

do I need to have an eigenvector for when lambda is 0?

OpenStudy (zarkon):

yes